# Extended AdS spacetime without boundaries and entanglement without holography

### v1

@@
Abstract
We glue together two copies of pure AdS spacetime along their conformal boundaries creating a manifold without boundaries. The resulting space, which in dimension $d+2$ we denote by $AdS^{d+2}_{\pm}$, has the topology of $S^{2}\times\Sigma^{d}$, where $\Sigma^{d}$ is a $d$-manifold without boundary. Acting with $\mathbb{Z}_{n}$ on the $S^{2}$ factor amounts to coupling a pair of membranes at the north and south poles of the 2-sphere. Moreover, extending the domain of the 2-sphere polar coordinate from $[0,\pi]$ to the interval $[0,(2N-1)\pi]$, where $N>1$, enables the coupling of one stack of $N$ coincident membranes at each pole of the 2-sphere ($2N$ membranes in total). Assuming the existence of a quantum gravity theory on the glued spacetime, we compute the classical approximation of the entanglement entropy across an entangling surface consisting of the two antipodal stacks of membranes. In odd dimensions, we find that each pair of membranes contributes with a quarter of their area to the entropy of the system. In even dimensions, the entropy is dependent on a boundary cutoff whose divergence can be suppressed by taking an infinite number of membranes. This cancellation yields exactly to a quarter of the area formula. The calculation does not require details of the quantum theory other than its infrared limit, that we assume to be Einstein gravity.
@@ 1 Introduction
Despite the central role of holography in crystallizing these ideas, many of the phenomena it captures appear to reflect more general principles. In particular, the appearance of area laws in vacuum entanglement entropy across a wide variety of quantum field theories [8, 9, 10, 11] suggests that the geometric nature of entanglement entropy may not require a dual boundary theory to be realized. Similarly, black hole entropy, with its universal structure, emerges from semiclassical gravity without invoking any particular holographic argument.
This raises the following question: to what extent the relation between entanglement entropy and the area of minimal surfaces can be understood and derived in non-holographic settings? In particular, can area laws of the Ryu–Takayanagi type be reproduced or reinterpreted intrinsically within the bulk spacetime without appealing to a dual boundary description?
@@
To this end, we construct a geometry consisting of two copies of AdS glued along their conformal boundaries. Unlike a single AdS copy, this doubled geometry allows for the coupling of an arbitrary number of membranes. These are minimal surfaces that appear as the fixed points of a bulk $\mathbb{Z}_{n}$ symmetry, and are organized into two antipodal stacks of $N$ coincident membranes each. We argue that the semiclassical approximation of the entanglement entropy between two spacetime subregions separated by the two stacks of membranes leads to a novel type of area law. This area law exhibits qualitatively different behavior depending on whether the spacetime dimension is even or odd.
In odd dimensions we find that each pair of membranes contributes with a factor of one-quarter of their area to the total entropy of the system; for the special case of a single pair of antipodal membranes, this entropy obeys the universal, one-quarter area law. In even dimension and for a finite number of membranes, the entropy depends on a boundary cutoff. However, this dependence can be suppressed by taking the limit of an infinite number membranes. In this limit, the resulting entanglement entropy obeys a Ryu–Takayanagi-like formula.
The article is organized as follows. In §2 we construct our extended geometry by gluing along the boundary two copies of AdS space. In §3 we show that the extended geometry naturally admits the coupling of an arbitrary number of membranes, organized into two antipodal stacks of $N$ coincident membranes each. In §4 we devise an spacetime subregion separated by the two stacks of membranes, and compute the associated entanglement entropy. We conclude in §5 with some final comments and remarks.
@@ 2 Extended AdS space without boundaries
To begin with, let us recall some elementary facts. Anti-de Sitter spacetime of dimension $d+2$ can be defined as an hyperboloid embedded in flat ambient space $\mathbb{R}^{2,d+1}$. Taking the ambient space coordinates to be $X^{A}=(T^{1},T^{2},X^{a})$, where $a=1,...,d+1$, the ambient space metric
@@ 2.1 Extended $AdS^{2}$
where $-\tfrac{\pi}{2}\leq\theta\leq\tfrac{\pi}{2}$ and $t\in\mathbb{R}$. It is direct to check that the choice above satisfies the hypersurface constraint (2.3). The corresponding induced metric on (a single copy of) $AdS^{2}$ is obtained by restricting the flat embedding space metric $\eta$ to the surface $X^{2}=-L^{2}$ by means of the the parametrization (2.5). This gives
@@
The advantage of these coordinates is that we can now glue two copies of $AdS^{2}$ simply by extending the range of the polar coordinate from $[-\frac{\pi}{2},\frac{\pi}{2}]$ to the interval $[-\frac{\pi}{2},\frac{3\pi}{2}]$. Indeed, taking these two copies to be defined as
$AdS^{2}_{+}$ :=$:=\bigg{\{}(T^{1},T^{2},X)\in\mathbb{R}^{2,1}|-\frac{\pi}{2}\leq\theta\leq\frac{\pi}{2}\bigg{\}}~{},$ (2.7)
$AdS^{2}_{-}$ :=$:=\bigg{\{}(T^{1},T^{2},X)\in\mathbb{R}^{2,1}|\quad\,\frac{\pi}{2}\leq\theta\leq\frac{3\pi}{2}\bigg{\}}~{},$ (2.8)
@@
can be defined as the result of overlapping the two boundaries by pairwise identifying the four points that compose them, that is, identifying the pair of points located at $\theta=\frac{\pi}{2}$ and the pair of points located at $\theta=-\frac{\pi}{2}$ and $\theta=\frac{3\pi}{2}$ Note that the range of the polar coordinate is chosen such that, in the resulting glued space, the north and south poles of the spatial circle are located at $\theta=0$ and $\pi$, respectively. The construction is depicted in figure 1.
@@ 2.2 Extended $AdS^{3}$
$AdS^{3}_{+}$ :=$:=\bigg{\{}(T^{1},T^{2},X^{1},X^{2})\in\mathbb{R}^{2,2}|~{}~{}0\leq\theta\leq\frac{\pi}{2}\bigg{\}}~{},$ (2.14)
$AdS^{3}_{-}$ :=$:=\bigg{\{}(T^{1},T^{2},X^{1},X^{2})\in\mathbb{R}^{2,2}|~{}~{}\frac{\pi}{2}\leq\theta\leq\pi\bigg{\}}~{},$ (2.15)
@@ 2.3 Extended $AdS^{d+2}$
and where $-\infty<(t^{1},t^{2},x^{i})<\infty$ are coordinates in the embedding subspace $\mathbb{R}^{2,d-1}\subset\mathbb{R}^{2,d+1}$. These are given by the restriction
@@
$AdS^{d+2}_{+}$ :=$:=\bigg{\{}(T^{1},T^{2},X^{i},X^{d},X^{d+1}\in\mathbb{R}^{2,d+1}~{}|~{}~{}0\leq\theta\leq\frac{\pi}{2}\bigg{\}}~{},$ (2.28)
$AdS^{d+2}_{-}$ :=$:=\bigg{\{}(T^{1},T^{2},X^{i},X^{d},X^{d+1}\in\mathbb{R}^{2,d+1}~{}|~{}~{}\frac{\pi}{2}\leq\theta\leq\pi\bigg{\}}~{},$ (2.29)
@@
$AdS^{d+2}_{\pm}=AdS^{d+2}_{+}\cup AdS^{d+2}_{-}=\bigg{\{}(T^{1},T^{2},X^{i},X^{d},X^{d+1}\in\mathbb{R}^{2,d+1}~{}|~{}~{}0\leq\theta\leq\pi\bigg{\}}~{}.$ (2.30)
@@ 3.1 Antipodal defects
The latter manifestly exhibits the presence conical defects at the north and south poles of the 2-sphere, both having a deficit angle given by $\Delta\phi=2\pi(1-\frac{1}{n})$. This maneuver amounts to thinking of the extended space $AdS^{d+2}_{\pm}$ as the smooth, $n\to 1$ limit of the product $S^{2}/\mathbb{Z}_{n}\times\Sigma^{d}$. As anticipated, for $n>1$ we denote the two singular surfaces supported by the antipodal defects by
@@ 3.2 Coupling an arbitrary number of membranes
where $y$ denotes the coordinates on $\Sigma_{+}$ and $\Sigma_{-}$. Note that since the extended manifold $AdS^{d+2}_{\pm}$ is the result of gluing two copies of AdS space along their boundaries, the corresponding boundary terms in (3.7) cancel each other due to their opposite orientations. Also note that, as a Riemannian submanifolds, $\Sigma_{+}$ and $\Sigma_{-}$ are both endowed with the same induced metric $h$, and the last two terms in (3.7) are simply the area functional of the singular surfaces coupled through the tension
@@
where $A(S)$ denotes the area of the hypersurface $S$. Evidently, $\theta$ can be extended as to run over the entire semi-positive real line $\theta\in[0,\infty]$, which translate into an infinite number of coupled membranes; this limit will be essential in regularizing the entropy divergence in even dimensions. A cartoon of the geometry with membranes is given in figure 3.
@@ 4 Entanglement entropy
Note that under this assumption and in this approximation, no details regarding the full quantum theory are required [^1] While the full quantum dynamics of gravity on manifolds without boundary remains largely unexplored, this assumption is common in studies of gravitational entropy and is consistent with the effective field theory approach to quantum gravity, where one expects universal IR behavior irrespective of UV completion [17]. The assumption has also proven to be useful in extracting reliable physical predictions at energy scales far below the Planck scale.
@@
for the Ricci scalar $R$ and cosmological constant $\Lambda$, both in dimension $d+2$, the on-shell value of the Euclidean [^3] The Euclidean action carries an overall minus sign compared to the Lorentzian functional (3.11). version of (3.11) is
$I[AdS_{\pm}/\mathbb{Z}_{n}]\approx\frac{d+1}{4nG}\bigg{[}\int_{0}^{\pi}\frac{\sin\theta}{(\cos\theta)^{d+2}}{\rm d}\theta\bigg{]}A(\Sigma)+\frac{N}{4G}\bigg{(}1-\frac{1}{n}\bigg{)}\Big{[}A(\Sigma_{+})+A(\Sigma_{-})\Big{]}~{},$ (4.10)
where $A(\Sigma)=\int{\rm d}y^{d}\sqrt{|{\rm det}\,h|}$, which due to the form of the metric (3.1) equals the areas $A(\Sigma_{+})$ and $A(\Sigma_{-})$.
@@
$\int_{0}^{\pi}\frac{\sin\theta}{(\cos\theta)^{d+2}}{\rm d}\theta=\int_{-1}^{-\varepsilon}\frac{{\rm d}u}{u^{d+2}}+\int_{\varepsilon}^{1}\frac{{\rm d}u}{u^{d+2}}=\begin{cases}\qquad~{}0&\text{for~{}~{}}d=2k+1\\ \frac{2}{d+1}(\frac{1}{\varepsilon^{d+1}}-1)&\text{for~{}~{}}d=2k\end{cases}$ (4.11)
where $k\geq 0$ and $\varepsilon$ is a boundary cutoff. Interestingly, for odd $d$ the integral above vanishes and so it does the cutoff dependence. For even $d$ the integral consists of a finite part plus a divergent term which in principle introduces a divergent term in the entanglement entropy. In what follows we analyze each of these cases separately.
4.1 Odd dimensions
The case of odd $d$ corresponds to odd-dimensional membranes $\Sigma^{d}_{+}$ and $\Sigma^{d}_{-}$ embedded in an odd-dimensional extended space $AdS^{d+2}_{\pm}$. As displayed in (4.11), in this case the singular integral vanishes and the on-shell action (4.10) becomes
$I[AdS_{\pm}/\mathbb{Z}_{n}]\approx\frac{N}{4G}\bigg{(}1-\frac{1}{n}\bigg{)}\Big{[}A(\Sigma_{+})+A(\Sigma_{-})\Big{]}~{}.$ (4.12)
Taking derivative with respect to $n$, the entanglement entropy (4.8) yields
$S_{\rm E}=N\,\bigg{[}\frac{A(\Sigma_{+}\cup\Sigma_{-})}{4G}\bigg{]}~{}.$ (4.13)
Note that for a single pair of membranes, that is when $N=1$, the formula above gives the one-quarter of the area law for the set of fixed points $\Sigma_{+}\cup\Sigma_{-}$ of the $\mathbb{Z}_{n}$ action. In general, when $N>1$, each pair of membranes contributes with one-quarter of their area.
4.2 Even dimensions
When the dimension of the membranes $\Sigma^{d}_{+}$ and $\Sigma^{d}_{-}$ is even, the integral (4.11) does not vanish and the on-shell action (4.10) becomes
@@ 4.2 Even dimensions
We observe that unlike the odd-dimensional case the entropy above depends on the boundary cutoff $\varepsilon$ in a way that resembles the usual leading, short-distance divergence of any entanglement entropy in dimension $d+2$, namely $1/\varepsilon^{d+1}$. However, in the present case, the freedom to couple an arbitrary number of membranes to the system amounts to take the double limit $N\to\infty$ and $\varepsilon\to 0$ in a controlled way such that the divergences $N-1/\varepsilon^{d+1}$ cancel each other. To this end, it suffices to take $N=10^{k}$ and $\varepsilon=10^{-k/(d+1)}$, in the limit $k\to\infty$.
It is important to point out that, according to (3.10), the $N\to\infty$ limit requires to extend the interval of the polar coordinate $\theta$ to the whole semi-positive real line. It is also crucial to observe that this cancelation of the divergences is only consistent in the tensionless limit $n\to 1$, in which the coupling constant $N(1-1/n)$ in the action (3.11) remains finite; when this is the case the entanglement entropy

### v2

@@
Extended AdS spacetime without boundaries
and entanglement without holography
Cesar Arias Departamento de Matemática, Pontificia Universidad Católica de Chile
Avenida Vicuña Mackenna 4860, Santiago, Chile
cesar.arias@uc.cl Abstract
We glue together two copies of pure AdS spacetime along their conformal boundaries creating a manifold without boundaries. The resulting space, which in dimension $d+2$ we denote by $AdS^{d+2}_{\pm}$, has the topology of $S^{2}\times\Sigma^{d}$, where $\Sigma^{d}$ is a $d$-manifold without boundary. Acting with $\mathbb{Z}_{n}$ on the $S^{2}$ factor amounts to coupling a pair of membranes at the north and south poles of the 2-sphere. Moreover, extending the domain of the 2-sphere polar coordinate from $0,\pi]$ to the interval $0,(2N-1)\pi]$, where $N>1$, enables the coupling of one stack of $N$ coincident membranes at each pole of the 2-sphere ($2N$ membranes in total). Assuming the existence of a quantum gravity theory on the glued spacetime, we compute the classical approximation of the entanglement entropy across an entangling surface consisting of the two antipodal stacks of membranes. We find that the resulting entropy exhibits a boundary cutoff divergence that can be canceled by taking the limit of an infinite number of membranes. This large-$N$ cancellation—possible only in the doubled, extended geometry without boundaries—yields a finite, universal quarter-area law. The calculation does not require details of the quantum theory other than its infrared limit, which we assume to be Einstein gravity.
@@ 1 Introduction
Despite the central role of holography in crystallizing these ideas, many of the phenomena it captures appear to reflect more general principles. In particular, the appearance of area laws in vacuum entanglement entropy across a wide variety of quantum systems [8, 9, 10, 11, 12, 13] suggests that the geometric nature of entanglement entropy may not require a dual boundary theory to be realized. Similarly, black hole entropy, with its universal structure, emerges from semiclassical gravity without invoking any particular holographic argument.
This raises the following question: to what extent can the relation between entanglement entropy and the area of minimal surfaces be understood and derived in non-holographic settings? In particular, can area laws of the Ryu–Takayanagi type be reproduced or reinterpreted intrinsically within the bulk spacetime without appealing to a dual boundary description?
@@
To this end, we construct a geometry consisting of two copies of AdS glued along their conformal boundaries. Unlike a single AdS copy, this doubled geometry allows for the coupling of an arbitrary number of membranes. These are minimal surfaces that appear as the fixed points of a bulk $\mathbb{Z}_{n}$ symmetry, and are organized into two antipodal stacks of $N$ coincident membranes each. We argue that, in a semiclassical approximation, the entanglement entropy between two spacetime subregions separated by the two stacks of membranes leads to a novel area law.
For a finite number $N$ of membranes, the entropy depends on a boundary cutoff and diverges as this cutoff is taken to zero. However, this divergence can be suppressed by taking the limit of an infinite number of membranes. In this limit, the resulting entanglement entropy becomes cutoff independent and obeys a Ryu–Takayanagi-like formula.
The article is organized as follows. In §2 we construct our extended geometry by gluing along the boundary two copies of AdS space. In §3 we show that the extended geometry naturally admits the coupling of an arbitrary number of membranes, organized into two antipodal stacks of $N$ coincident membranes each. In §4 we devise a spacetime subregion separated by the two stacks of membranes, and compute the associated entanglement entropy. We conclude in §5 with some final comments and remarks.
@@ 2 Extended AdS space without boundaries
To begin with, let us recall some elementary facts. Anti-de Sitter spacetime of dimension $d+2$ can be defined as a hyperboloid embedded in flat ambient space $\mathbb{R}^{2,d+1}$. Taking the ambient space coordinates to be $X^{A}=(T^{1},T^{2},X^{a})$, where a=1,,d+1a=1,...,d+1, the ambient space metric
@@ 2.1 Extended $AdS^{2}$
where $\tfrac{\pi}{2}\leq\theta\leq\tfrac{\pi}{2}$ and $t\in\mathbb{R}$. It is direct to check that the choice above satisfies the hypersurface constraint (2.3). The corresponding induced metric on (a single copy of) $AdS^{2}$ is obtained by restricting the flat embedding space metric $\eta$ to the surface $X^{2}=-L^{2}$ by means of the the parametrization (2.5). This gives
@@
The advantage of these coordinates is that we can now glue two copies of $AdS^{2}$ simply by extending the range of the polar coordinate from $\frac{\pi}{2},\frac{\pi}{2}]$ to the interval $\frac{\pi}{2},\frac{3\pi}{2}]$. Indeed, taking these two copies to be defined as
$AdS^{2}_{+}$ $:=\bigg\{(T^{1},T^{2},X)\in\mathbb{R}^{2,1}|-\frac{\pi}{2}\leq\theta\leq\frac{\pi}{2}\bigg\}~,$ (2.7)
$AdS^{2}_{-}$ $:=\bigg\{(T^{1},T^{2},X)\in\mathbb{R}^{2,1}|\quad\,\frac{\pi}{2}\leq\theta\leq\frac{3\pi}{2}\bigg\}~,$ (2.8)
@@
can be defined as the result of overlapping the two boundaries by pairwise identifying the four points that compose them, that is, identifying the pair of points located at $\theta=\frac{\pi}{2}$ and the pair of points located at $\theta=-\frac{\pi}{2}$ and $\theta=\frac{3\pi}{2}$ . Note that the range of the polar coordinate is chosen such that, in the resulting glued space, the north and south poles of the spatial circle are located at $\theta=0$ and $\pi$, respectively. The construction is depicted in figure 1.
@@ 2.2 Extended $AdS^{3}$
$AdS^{3}_{+}$ $:=\bigg\{(T^{1},T^{2},X^{1},X^{2})\in\mathbb{R}^{2,2}|~~0\leq\theta\leq\frac{\pi}{2}\bigg\}~,$ (2.14)
$AdS^{3}_{-}$ $:=\bigg\{(T^{1},T^{2},X^{1},X^{2})\in\mathbb{R}^{2,2}|~~\frac{\pi}{2}\leq\theta\leq\pi\bigg\}~,$ (2.15)
@@ 2.3 Extended $AdS^{d+2}$
and where $\infty<(t^{1},t^{2},x^{i})<\infty$ are coordinates in the embedding subspace $\mathbb{R}^{2,d-1}\subset\mathbb{R}^{2,d+1}$. These are given by the restriction
@@
$AdS^{d+2}_{+}$ $:=\bigg\{(T^{1},T^{2},X^{i},X^{d},X^{d+1})\in\mathbb{R}^{2,d+1}~|~~0\leq\theta\leq\frac{\pi}{2}\bigg\}~,$ (2.28)
$AdS^{d+2}_{-}$ $:=\bigg\{(T^{1},T^{2},X^{i},X^{d},X^{d+1})\in\mathbb{R}^{2,d+1}~|~~\frac{\pi}{2}\leq\theta\leq\pi\bigg\}~,$ (2.29)
@@
$AdS^{d+2}_{\pm}=AdS^{d+2}_{+}\cup AdS^{d+2}_{-}=\bigg\{(T^{1},T^{2},X^{i},X^{d},X^{d+1})\in\mathbb{R}^{2,d+1}~|~~0\leq\theta\leq\pi\bigg\}~.$ (2.30)
@@ 3.1 Antipodal defects
The latter manifestly exhibits the presence of conical defects at the north and south poles of the 2-sphere, both having a deficit angle given by $\Delta\phi=2\pi(1-\frac{1}{n})$. This maneuver amounts to thinking of the extended space $AdS^{d+2}_{\pm}$ as the smooth, $n\to 1$ limit of the product $S^{2}/\mathbb{Z}_{n}\times\Sigma^{d}$. As anticipated, for $n>1$ we denote the two singular surfaces supported by the antipodal defects by
@@ 3.2 Coupling an arbitrary number of membranes
where $y$ denotes the coordinates on $\Sigma_{+}$ and $\Sigma_{-}$. Note that since the extended manifold $AdS^{d+2}_{\pm}$ is the result of gluing two copies of AdS space along their boundaries, the corresponding boundary terms in (3.7) cancel each other due to their opposite orientations. Also note that, as Riemannian submanifolds, $\Sigma_{+}$ and $\Sigma_{-}$ are both endowed with the same induced metric $h$, and the last two terms in (3.7) are simply the area functional of the singular surfaces coupled through the tension
@@
where $A(S)$ denotes the area of the hypersurface $S$. Evidently, $\theta$ can be extended as to run over the entire semi-positive real line $\theta\in[0,\infty]$, which translates into an infinite number of coupled membranes; this limit will be essential in regularizing the entropy divergence in even dimensions. A cartoon of the geometry with membranes is given in figure 3.
@@ 4 Entanglement entropy
Note that under this assumption and in this approximation, no details regarding the full quantum theory are required [^1] While the full quantum dynamics of gravity on manifolds without boundary remains largely unexplored, this assumption is common in studies of gravitational entropy and is consistent with the effective field theory approach to quantum gravity, where one expects universal IR behavior irrespective of UV completion [19]. The assumption has also proven to be useful in extracting reliable physical predictions at energy scales far below the Planck scale..
@@
for the Ricci scalar $R$ and cosmological constant $\Lambda$, both in dimension $d+2$, the Euclidean [^3] The Euclidean action carries an overall minus sign compared to the Lorentzian functional (3.11). on-shell value of (3.11) is
$I[AdS_{\pm}/\mathbb{Z}_{n}]\approx\frac{d+1}{4nG}\bigg[\int_{0}^{\pi}\frac{\sin\theta}{|\cos\theta|^{d+2}}{\rm d}\theta\bigg]A(\Sigma)+\frac{N}{4G}\bigg(1-\frac{1}{n}\bigg)\Big[A(\Sigma_{+})+A(\Sigma_{-})\Big]~.$ (4.10)
In the previous expression we have denoted by $A(\Sigma)=\int{\rm d}y^{d}\sqrt{|{\rm det}\,h|}$, which, due to the form of the metric (3.1) equals the areas $A(\Sigma_{+})$ and $A(\Sigma_{-})$.
@@
$\int_{0}^{\pi}\frac{\sin\theta}{|\cos\theta|^{d+2}}{\rm d}\theta=\int_{-1}^{-\varepsilon}\frac{{\rm d}u}{|u|^{d+2}}+\int_{\varepsilon}^{1}\frac{{\rm d}u}{|u|^{d+2}}=\frac{2}{d+1}\bigg(\frac{1}{\varepsilon^{d+1}}-1\bigg)$ (4.11)
where $\varepsilon>0$ is a boundary cutoff.
By virtue of the integral (4.11), the on-shell action (4.10) becomes
@@
We observe that the entropy above depends on the boundary cutoff $\varepsilon$ in a way that resembles the usual leading, short-distance divergence of any entanglement entropy in dimension $d+2$, namely $1/\varepsilon^{d+1}$. However, in the present case, the freedom to couple an arbitrary number of membranes to the system amounts to take the double limit $N\to\infty$ and $\varepsilon\to 0$ in a controlled way such that the divergences $N-1/\varepsilon^{d+1}$ cancel each other. To this end, it suffices to take $N=10^{k}$ and $\varepsilon=10^{-k/(d+1)}$, with $k\to\infty$.
It is important to point out that, according to (3.10), the $N\to\infty$ limit requires to extend the interval of the polar coordinate $\theta$ to the whole semi-positive real line. It is also crucial to observe that this cancelation of the divergences is only consistent in the tensionless limit $n\to 1$, in which the coupling constant N(11/n)N(1-1/n) in the action (3.11) remains finite; when this is the case the entanglement entropy